OpenStudy (maddy1251):

Quick question regarding Logs. Last one tonight I promise! I have to decide which of the 2 are correct. I believe it to be the first one, but I am unsure. Just to have someone reassure me would mean a lot!! My two choices are logM^p=p x logM or (logM)^p= p x logM Logarithm power rule logb(x y) = y × logb(x) this is the rule I am going by.

Nnesha (nnesha):

power rule \[\large\rm log_b x^y = y \log_b x\]

Nnesha (nnesha):

what's ur question ??

OpenStudy (maddy1251):

Which of the following statements are true? <---

Nnesha (nnesha):

ohh i see so that's the power rule ive posted above which can be written as \[\large\rm log_b (x)^y = y \log_b x\]

OpenStudy (maddy1251):

My one choice I have selected as being true is logM + logN = log(MN)

Nnesha (nnesha):

that's correct and that's the product rule!

OpenStudy (maddy1251):

I figured that was the rule they were looking for. Wouldn't that make it logM^p=p x logM.

OpenStudy (maddy1251):

I know either of the two logs I posted is true, I just don't know which one it is. :P

Nnesha (nnesha):

by this `x` do you mean multiplication sign or x variable ? just making sure

OpenStudy (maddy1251):

multiplication. My bad

Nnesha (nnesha):

good that's the correct `power rule `

OpenStudy (maddy1251):

Okay. For the second one I posted (logM)^p= p x logM, what makes it different?

Nnesha (nnesha):

\[\huge\rm log~ m^\color{Red}{p} = \color{red}{p}* log~m\]

Nnesha (nnesha):

i don't see any difference let say p =2 and m =3\[\log (3)^2 = (\log(3))^2\]

OpenStudy (maddy1251):

It's just the use of parentheses that changes that?

OpenStudy (maddy1251):

Is it basically like a trick to make you think it's not a choice?

Nnesha (nnesha):

well its gonna move to the front anyway i mean we have to apply the power rule \[\rm (\log m)^p = p \log m\]

OpenStudy (triciaal):

if I am reading it right yes the first one is correct by power rule the difference that I see is the log of the number then to the power or the log of of the number raised to the power

Nnesha (nnesha):

logb(x y) = y × logb(x) this is not correct \[\rm y *\log_b (x) = \log (x*y)\]

Nnesha (nnesha):

\[\log_b x+\log_b y = \log(x*y)\]

OpenStudy (triciaal):